# Multi-scale analysis of flow over heterogeneous urban environments

### v1

@@ 5 Conclusions
This paper presented a multi-scale framework that permits systematic investigation of flows over heterogeneous urban areas. The use of convolution filters brings the Whitaker volume-averaging approach (Whitaker 1999) in line with the theory of large-eddy simulation, and permits for a computationally efficient procedure to carry out coarse-graining, which is particularly important for analysis of high-resolution simulations as considered here.
@@
One of the central questions the paper aimed to address was: ”What are the requirements for non-building-resolving Numerical-Weather-Prediction models (NWP) as their resolution is increased?” To answer this question, a quasi-realistic case was considered inspired by the building height distribution of London. The multi-scale framework revealed that the flow has a characteristic length scale $\ell\approx 50$ m, which is the resolution at which as much variance is resolved as is unresolved. This corresponds to roughly twice the building width. Heterogeneity effects, which challenge the NWP parameterisations for the land surface, become significant at resolutions of about 200 m for the case under consideration.
@@
Given that future NWP models will likely run at resolutions at which heterogeneity effects are important, an immediate follow-up question is how to incorporate heterogeneity. This requires a land-surface model that resolves the vertical distribution of the flow and turbulence (Lean et al. 2024) and requires: i) a model for the distribution of the drag over the canopy; and ii) a model for the unresolved turbulence and dispersive stresses in the canopy. Since NWP models do not resolve the urban morphology and it is undesirable to have to modify the governing equations inside the canopy (e.g. the mass conservation equation), superficial (comprehensive) averages are the appropriate formulation to consider, see (17). It is therefore highly desirable that the drag distribution and unresolved stress parameterisation are formulated in terms of superficial quantities also.
@@
It was shown that the distributed drag parameterisation of Sützl et al. (2020), which was obtained from plane-averaged analysis of an ensemble of eight different urban landscapes, can be successfully extended to a local formulation, cf (40). The parameterisation generalises excellently, with an $R^{2}$ value of 0.929 at 96 m resolution. The attractive feature of this parameterisation is that it only requires the local distributed frontal area $\zeta_{L}$ which can be calculated a priori and can be straightforwardly incorporated in NWP models (Sützl et al. 2021). Although this parameterisation is excellent for predicting the local drag distribution, the actual drag requires knowledge of the local surface stress which remains challenging and requires further work; a local version of the Coceal and Belcher (2004); Belcher (2005) parameterisation (43) has a normalised root-mean-square error is 71% at 100 m resolution.
@@
Several closures were considered for the unresolved and turbulent stresses in the urban canopy layer. The two best-performing models were the $k-\omega$ model, which requires information about the turbulence kinetic energy $k$ and the inverse turbulence timescale $\omega$, and a new stress parameterisation based on the distributed drag, cf (46), that does not require further parameters. At a resolution of 96 m, the latter performs better than the former, although both parameterisations have a substantial scatter associated with local heterogeneity effects.
@@
The results reported here indicate that it is possible to incorporate substantial heterogeneity effects in non-building-resolving NWP models with relative ease. However, further work is required. First, it is desirable to ensure the validity of the parameterisations (36) and (46) on a much larger flow database, e.g. that of Lu et al. (2024). Furthermore, the atmosphere is non-neutral most of the time, and distributed versions of the sensible and latent heat-fluxes, including the interaction with the surface stress and needs to be explored.
@@ Acknowledgements
The support of the ARCHER2 UK National Supercomputing Service (project ARCHER2-eCSE05-3) and the NESRC highlight grant ASSURE: Across-Scale ProcesseS in Urban Environment (NE/W002868/1) is acknowledged.
@@ Appendix A Volume-averaging with convolution filters
where $\boldsymbol{x}$ is the 3-D location and $\int\mathcal{G}(\boldsymbol{x})\textnormal{d}\boldsymbol{x}=1$, i.e. the kernel is normalised. Note that in (47), integration only takes place over the fluid phase. The fluid volume fraction $\varepsilon=V_{f}/V$, where $V_{f}$ is the volume occupied by fluid inside the averaging volume, is given by
@@ A.1 Volume to area integration
$\mathcal{G}(\boldsymbol{x})=\mathcal{A}(\boldsymbol{x}_{\perp})\delta(z)\,,$ (49)
where $\delta$ is the Kronecker delta and $\boldsymbol{x}=[\boldsymbol{x}_{\perp},z]^{T}$. With this definition of $\mathcal{G}$, the volume integral will become an area integral. The planar filter is also normalised:
$\int\mathcal{G}(\boldsymbol{x})\textnormal{d}\boldsymbol{x}=\int\int\mathcal{A}(\boldsymbol{x}_{\perp})\delta(z)\textnormal{d}\boldsymbol{x}_{\perp}\textnormal{d}z=\int\mathcal{A}(\boldsymbol{x}_{\perp})\textnormal{d}\boldsymbol{x}_{\perp}=1\,.$
Using (49), the superficial average (47) becomes
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\varphi(\boldsymbol{y}_{\perp},z)\textnormal{d}\boldsymbol{y}_{\perp}\,.$ (50)
@@
$\varepsilon(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}\,,$ (51)
@@ A.2 Differentiation rules
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\int_{\Omega_{f}}\mathcal{A}\nabla\cdot\boldsymbol{F}\textnormal{d}\boldsymbol{y}_{\perp}=\int_{\Omega_{f}}\nabla\cdot(\mathcal{A}\boldsymbol{F})\textnormal{d}\boldsymbol{y}_{\perp}-\int_{\Omega_{f}}\boldsymbol{F}\cdot\nabla\mathcal{A}\textnormal{d}\boldsymbol{y}_{\perp}$ (52)
@@
$\int_{\Omega_{f}}\nabla\cdot\boldsymbol{G}\textnormal{d}\boldsymbol{y}_{\perp}=\frac{\partial}{\partial z}\int_{\Omega_{f}}G_{z}\textnormal{d}\boldsymbol{y}_{\perp}-\oint_{\partial\Omega_{f}}\frac{\boldsymbol{G}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s.$ (53)
Using (53), (52) becomes
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\frac{\partial\left\langle{\smash{F_{z}}}\right\rangle}{\partial z}-\int_{\Omega_{f}}\boldsymbol{F}_{\perp}\cdot\nabla_{\perp}\mathcal{A}\textnormal{d}\boldsymbol{y}_{\perp}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{\boldsymbol{F}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s,$ (54)
@@
$\begin{split}\int_{\Omega_{f}}\boldsymbol{F}_{\perp}(\boldsymbol{y}_{\perp})\cdot\nabla_{\perp}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}&=\int_{\Omega_{f}}\boldsymbol{F}_{\perp}(\boldsymbol{y}_{\perp})\cdot\frac{\partial}{\partial\boldsymbol{y}_{\perp}}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}\\ &=-\int_{\Omega_{f}}\boldsymbol{F}_{\perp}(\boldsymbol{y}_{\perp})\cdot\frac{\partial}{\partial\boldsymbol{x}_{\perp}}\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}\\ &=-\frac{\partial}{\partial\boldsymbol{x}_{\perp}}\cdot\int_{\Omega_{f}}\boldsymbol{F}_{\perp}(\boldsymbol{y}_{\perp})\mathcal{A}(\boldsymbol{x}_{\perp}-\boldsymbol{y}_{\perp})\textnormal{d}\boldsymbol{y}_{\perp}\\ &=-\nabla_{\perp}\cdot\left\langle{\smash{\boldsymbol{F}_{\perp}}}\right\rangle,\end{split}$
@@
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\nabla\cdot\left\langle{\smash{\boldsymbol{F}}}\right\rangle-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{\boldsymbol{F}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s.$ (55)
Substitution of $\boldsymbol{F}=\varphi\boldsymbol{e}_{i}$ where $i\in\{x,y,z\}$ into (55) results in the the spatial averaging theorem for scalars:
$\left\langle{\smash{\nabla\varphi}}\right\rangle=\nabla\left\langle{\smash{\varphi}}\right\rangle-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{\boldsymbol{N}}{|\boldsymbol{N}_{\perp}|}\textnormal{d}s\,.$ (56)

### v2

@@ 5 Conclusions
A multi-scale framework was developed for investigating the flows over heterogeneous urban areas. At the core of the framework is a coarse-graining method that employs convolution filters to obtain coarse-grained fields from a building-resolved high-resolution simulation. By employing Fast-Fourier transforms, a computationally efficient procedure was developed to carry out coarse-graining, which is particularly important for the analysis of high-resolution simulations as considered here.
@@
One of the central questions the paper aimed to address was: ”What are the requirements for non-building-resolving NWP models as their resolution is increased?” To answer this question, a quasi-realistic case was considered, inspired by the building height distribution of London. The multi-scale framework revealed that the flow has a characteristic length scale $\ell\approx 50$ m, which is the resolution at which as much variance is resolved as is unresolved. This corresponds to roughly twice the building width. Heterogeneity effects, which challenge the NWP parameterisations for the land surface, become significant at resolutions of about 200 m and below for the case under consideration.
@@
Given that future NWP models will likely run at resolutions at which heterogeneity effects are important, an immediate follow-up question is how to incorporate heterogeneity. This requires a land-surface model that resolves the vertical distribution of the flow and turbulence (Lean et al. 2024) and requires: i) a model for the distribution of the drag over the canopy; and ii) a model for the unresolved turbulence and dispersive stresses in the canopy. Since NWP models do not resolve the urban morphology and it is undesirable to have to modify the governing equations inside the canopy (e.g. the mass conservation equation), superficial (comprehensive) averages are the appropriate formulation to consider, see Eq. (17). It is therefore highly desirable that the drag distribution and unresolved stress parameterisation are formulated in terms of superficial quantities also.
@@
It was shown that the distributed drag parameterisation of Sützl et al. (2020), which was obtained from plane-averaged analysis of an ensemble of eight different urban landscapes, can be successfully extended to a local formulation, cf Eq. (40). The parameterisation generalises excellently, with an $R^{2}$ value of 0.929 at 96 m resolution. The attractive feature of this parameterisation is that it only requires the local distributed frontal area $\zeta_{L}$ which can be calculated a priori and can be straightforwardly incorporated in NWP models (Sützl et al. 2021). Although this parameterisation is excellent for predicting the local drag distribution, the actual drag requires knowledge of the local surface stress which remains challenging and requires further work; a local version of the Coceal and Belcher (2004); Belcher (2005) parameterisation Eq. (43) has a normalised root-mean-square error of 71% at 100 m resolution.
@@
Several closures were considered for the unresolved and turbulent stresses in the urban canopy layer. The two best-performing models were the $k-\omega$ model, which requires information about the turbulence kinetic energy $k$ and the inverse turbulence timescale $\omega$, and a new stress parameterisation based on the distributed drag, cf Eq. (46), that does not require further parameters. At a resolution of 96 m, the latter performs better than the former, although both parameterisations have a substantial scatter associated with local heterogeneity effects.
@@
The results reported here indicate that it is possible to incorporate substantial heterogeneity effects in non-building-resolving NWP models with relative ease. However, further work is required. First, it is desirable to ensure the validity of the parameterisations Eq. (36) and Eq. (46) on a much larger flow database, e.g. that of Lu et al. (2024). Furthermore, the atmosphere is non-neutral most of the time, and distributed versions of the sensible and latent heat fluxes, including the interaction with the surface stress, need to be explored.
@@ Acknowledgements
The support of the ARCHER2 UK National Supercomputing Service (project ARCHER2-eCSE05-3) and the NERC highlight grant ASSURE: Across-Scale ProcesseS in Urban Environment (NE/W002868/1) is acknowledged.
Data availability
The data that support the findings of this study, as well as the coarsegraining routines are openly available in Zenodo at Van Reeuwijk and Huang (2025).
@@ Appendix A Volume-averaging with convolution filters
where $\boldsymbol{x}$ is the 3-D location and $\int\mathcal{G}(\boldsymbol{x})\textnormal{d}\boldsymbol{x}=1$, i.e. the kernel is normalised. Note that in Eq. (47), integration only takes place over the fluid phase. The fluid volume fraction $\varepsilon=V_{f}/V$, where $V_{f}$ is the volume occupied by fluid inside the averaging volume, is given by
@@ A.1 Volume to area integration
$\mathcal{G}(\boldsymbol{x})=\mathcal{A}(\boldsymbol{x}_{h})\delta(z)\,$, (49)
where $\delta$ is the Kronecker delta and $\boldsymbol{x}=[\boldsymbol{x}_{h},z]^{T}$. With this definition of $\mathcal{G}$, the volume integral will become an area integral. The planar filter is also normalised:
$\int\mathcal{G}(\boldsymbol{x})\textnormal{d}\boldsymbol{x}=\int\int\mathcal{A}(\boldsymbol{x}_{h})\delta(z)\textnormal{d}\boldsymbol{x}_{h}\textnormal{d}z=\int\mathcal{A}(\boldsymbol{x}_{h})\textnormal{d}\boldsymbol{x}_{h}=1\,.$
Using Eq. (49), the superficial average Eq. (47) becomes
$\left\langle{\smash{\varphi}}\right\rangle(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\varphi(\boldsymbol{y}_{h},z)\textnormal{d}\boldsymbol{y}_{h}\,.$ (50)
@@
$\varepsilon(\boldsymbol{x})=\int_{\Omega_{f}(\boldsymbol{x})}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}\,$, (51)
@@ A.2 Differentiation rules
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\int_{\Omega_{f}}\mathcal{A}\nabla\cdot\boldsymbol{F}\textnormal{d}\boldsymbol{y}_{h}=\int_{\Omega_{f}}\nabla\cdot(\mathcal{A}\boldsymbol{F})\textnormal{d}\boldsymbol{y}_{h}-\int_{\Omega_{f}}\boldsymbol{F}\cdot\nabla\mathcal{A}\textnormal{d}\boldsymbol{y}_{h}$ (52)
@@
$\int_{\Omega_{f}}\nabla\cdot\boldsymbol{G}\textnormal{d}\boldsymbol{y}_{h}=\frac{\partial}{\partial z}\int_{\Omega_{f}}G_{z}\textnormal{d}\boldsymbol{y}_{h}-\oint_{\partial\Omega_{f}}\frac{\boldsymbol{G}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{h}|}\textnormal{d}s.$ (53)
Using Eq. (53), Eq. (52) becomes
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\frac{\partial\left\langle{\smash{F_{z}}}\right\rangle}{\partial z}-\int_{\Omega_{f}}\boldsymbol{F}_{h}\cdot\nabla_{h}\mathcal{A}\textnormal{d}\boldsymbol{y}_{h}-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{\boldsymbol{F}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{h}|}\textnormal{d}s,$ (54)
@@
$\begin{split}\int_{\Omega_{f}}\boldsymbol{F}_{h}(\boldsymbol{y}_{h})\cdot\nabla_{h}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}&=\int_{\Omega_{f}}\boldsymbol{F}_{h}(\boldsymbol{y}_{h})\cdot\frac{\partial}{\partial\boldsymbol{y}_{h}}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}\\ &=-\int_{\Omega_{f}}\boldsymbol{F}_{h}(\boldsymbol{y}_{h})\cdot\frac{\partial}{\partial\boldsymbol{x}_{h}}\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}\\ &=-\frac{\partial}{\partial\boldsymbol{x}_{h}}\cdot\int_{\Omega_{f}}\boldsymbol{F}_{h}(\boldsymbol{y}_{h})\mathcal{A}(\boldsymbol{x}_{h}-\boldsymbol{y}_{h})\textnormal{d}\boldsymbol{y}_{h}\\ &=-\nabla_{h}\cdot\left\langle{\smash{\boldsymbol{F}_{h}}}\right\rangle,\end{split}$
@@
$\left\langle{\smash{\nabla\cdot\boldsymbol{F}}}\right\rangle=\nabla\cdot\left\langle{\smash{\boldsymbol{F}}}\right\rangle-\oint_{\partial\Omega_{f}}\mathcal{A}\frac{\boldsymbol{F}\cdot\boldsymbol{N}}{|\boldsymbol{N}_{h}|}\textnormal{d}s.$ (55)
Substitution of $\boldsymbol{F}=\varphi\boldsymbol{e}_{i}$ where $i\in\{x,y,z\}$ into Eq. (55) results in the the spatial averaging theorem for scalars:
$\left\langle{\smash{\nabla\varphi}}\right\rangle=\nabla\left\langle{\smash{\varphi}}\right\rangle-\oint_{\partial\Omega_{f}}\mathcal{A}\varphi\frac{\boldsymbol{N}}{|\boldsymbol{N}_{h}|}\textnormal{d}s\,.$ (56)